Two flat panes of glass (80 cm long, 15 cm wide, 4
Answer #1

2t = 2\theta= m\lambda   for a dark fringe

2t = 2\theta= (2m+1)\lambda/2   for a bright fringe

\Delta x=L\frac{\lambda }{2t}

\Delta x=2.5cm=2.5\times 10^{-2}m

L=80cm=80\times 10^{-2}m

\lambda =650nm=650\times 10^{-9}m

\Delta x=L\frac{\lambda }{2t}\Rightarrow t=\frac{L\lambda }{2\Delta x}

t=\frac{L\lambda }{2\Delta x}

t=\frac{80\times 10^{-2}m\times 650\times 10^{-9}m}{2\times 2.5\times 10^{-2}m}

t=\frac{52\times 10^{-8}m^2}{5\times 10^{-2}m}=10.4\times 10^{-6}m

t=10.4\times \mu m